OpenStudy (anonymous):

How long does it take (in minutes) for light to reach Jupiter from the Sun, a distance of 8.390 × 108 km?

OpenStudy (compphysgeek):

the speed of light is about \(c = 3 \times 10^8 m/s\) or \(c = 3 \times 10^5 km/s\). multiply c by \(60 s/min\) and you'll get \(1.80 \times 10^7 km/min\). Now just divide the distance by the speed and you'll have the time.